# P-NAL: an Effective and Interpretable Entity Alignment Method

### v1

@@ 3.3. Utilizing Literal Values
This path seems straightforward, however we can have a deeper understanding of it. Language models used for the embedding process of EA are distinct information sources other than the KG itself. The deep language model which has the ability of aligning or translating entity names can be seen as a generalized alignment model that aligns morphemes, words, entities and concepts. The pretraining corpus of it consists of sentences, although the sentences do not possesses explicit structures, they can be understood by the model by transforming them into complex logical forms. However, such transformation (if exist) and the logical forms are implicitly expressed in the model parameters. To summarize, our similarity inference’s type II path can be seen as the aggregation of multiple (virtual) complex logical paths. The aggregation result is represented into the vector space by the language model.
Attribute value feature’s cosine similarity is used to convert to the truth-value of premise statement (5) if $x_{1}$ and $x_{2}$ are attribute values. The corresponding truth-value is
$\left\langle f=sim(x_{1},x_{2}),\ c=\-sim(x_{1},x_{2})\right\rangle$
The idea is that the deep learning model’s result which has higher similarity is usually more verifiable. There are thousands of distinct attribute values in a KG, so for an attribute value we only consider the $K_{value}$ most similar (but not identical) values in the other KG to prevent an explosive number of value similarities. $K_{value}$ is a hyper-parameter and in implementation we set $K_{value}$ to 1.
3.4. Matching Algorithm
@@
4       recursively_delete($e_{1},null$);
@@
/* $e_{1}$ is the entity to be processed and we assume that $e_{1}$ belongs to the left graph, similarly otherwise. Entity $e_{prev}$ represents the previous entity, that is the processed entity of the recursion parent. */
output : entity $e_{return}$ which represents the final alignment for $e_{1}$
@@ 3       $e_{2}$leftarrow $predicate\_term$ of $sentence$;
5             $e_{return}$leftarrow $e_{prev}$;
@@
9             $e_{3}$leftarrow recursively_delete($e_{2},e_{1}$);
@@ 11                   $e_{return}$leftarrow $e_{2}$;
Algorithm 2 recursively delete
There are 1-to-1 assumptions in some EA datasets (such as $DBP\-15K$) and it is a useful information for alignment. Formally, we define the 1-to-1 assumption as follows: first, there is a range of alignable entities $A_{1}\subset\mathcal{E}_{1}$ and $A_{2}\subset\mathcal{E}_{2}$ (for $DBP15K$, $A_{1}\subsetneqq\mathcal{E}_{1}$). Second, the equivalence between $A_{1}$ and $A_{2}$ is a bijection. Note that the assumption does not have aligning regularity for entities outside the range except that they can’t be aligned with entities inside the range. For example, $DBP15K_{ZH\_EN}$’s $ZH$ graph has a 15,000 sized entity set $A_{1}$ and $EN$ graph has a 15,000 sized entity set $A_{2}$. It is informed that every entity in $A_{1}$ must have a unique entity in $A_{2}$ as its alignment counterpart. Many ranking-based EA methods leverages the 1-to-1 range assumption, however, PARIS do not. Therefore, in implementation in order to leverage the range assumption we take the set $A_{1}$ and $A_{2}$ as input and filters out any alignment sentence that aligns $A_{1}$ to $\mathcal{E}_{2}\setminus A_{2}$ or $\mathcal{E}_{1}\setminus A_{1}$ to $A_{2}$.
Align-bridge’s similarities (type I path) are naturally sparse, because it only considers the entity pairs which is effectively linked by the logical path. Entity name/description’s similarities (type II path) are dense, however, it is noisy and most of the similarities are useless. P-NAL’s overall algorithm (depicted in the next Section) exhaustively search for and stores the two types of similarity sentences for a specific to-be-aligned entity versus any entity in the other KG. Then, because of the sparsity of informative similarity signal, the similarity sentences is rearranged into ordered linked list, one list for a specific to-be-aligned entity. The sentences are ordered (descending) by its $expectation$ value. We only store the top $K_{sim}$ similarity sentences in the linked list, where $K_{sim}$ is a model hyper-parameter.
Then we perform a recursive bidirectional matching algorithm (rBMat) which has similar idea with BMat  (Tang et al., 2020) but different implementation. See Algorithm 1 and Algorithm 2 for details. The main idea is to recursively delete the similarity sentences that don’t conform the 1-to-1 assumption. Considering sorting cost, our rBMat has $O(kn^{2})$ time complexity and $O(kn)$ space complexity, where $k$ represents $K_{sim}$ and $k\ll n$.
We found that there are still some mismatches after performing rBMat algorithm and most of them share a same pattern. For example, $e_{1}\leftrightarrow e_{2}$ and $e_{3}\leftrightarrow e_{4}$ are two ground truth pairs, however, rBMat’s result is $e_{1}\leftrightarrow e_{3}$ and $e_{2}\leftrightarrow e_{4}$. We implements a simple “modify matches” algorithm to handle this. We exhaustively search for and modify (swap the alignment) the cases in which $expt(e_{1}\leftrightarrow e_{2})+expt(e_{3}\leftrightarrow e_{4})$ is greater than $expt(e_{1}\leftrightarrow e_{3})+expt(e_{2}\leftrightarrow e_{4})$, where $expt$ is the $expectation$ of the truth-value.
3.5. Overall structure of P-NAL
Generally speaking, our method adopts the same optimization method and iteration strategy as PARIS’s implementation, with some minor differences. The overall structure of supervised P-NAL algorithm is elaborated in Algorithm 3. Line 1-3 finetunes the BERT unit. Line 6-13 performs similarity inference and relational alignment. Line 14-19 performs the proposed matching algorithm to obtain alignment results of an iteration. $End\_iteration$ is a hyper-parameter. Note that the inference within each iteration benefits from the alignment results (both entities and relations) of the previous iteration. Registering evidential information (line 10 in Algorithm 3) means memorizing which premises form the specific align-bridge and such information will be used to generate evidence log file.
@@ 3.5. Overall structure of P-NAL
8             for $x_{1}$, $x_{2}$, $y_{2}$ that forms a sound align-bridge path with $y_{1}$ do
                   /* Type I path. The paths are searched in a depth-first manner */
9                   update the estimations of $w$ and $w+$ for relational inheritance;
10                   perform inference with the inference steps and inference rules in Section 3.1;
11                   register evidential information of the path;
12
@@
15                   integrate the similarity with prior conclusions of align-bridge by revision rule;
16

### v2

@@ D.2 Understanding of type II path
Type II path seems straightforward, however we can have a deeper understanding of it. Language models used for the embedding process of EA are distinct information sources other than the KG itself. The deep language model which has the ability of aligning or translating entity names can be seen as a generalized alignment model that aligns morphemes, words, entities and concepts. The pretraining corpus of it consists of sentences, although the sentences do not possesses explicit structures, they can be understood or parsed by the model by transforming them into complex logical forms. However, such transformation (if exist) and the logical forms are implicitly expressed in the model parameters and intermediate layer vector representations. To summarize, our similarity inference’s type II path can be seen as the aggregation of multiple virtual complex logical paths. The aggregation result is represented into the vector space by the language model.
D.3 1-to-1 Assumption
There are 1-to-1 assumptions in some EA datasets (such as $DBP\-15K$) and it is a useful information for alignment. Formally, we define the 1-to-1 assumption as follows: first, there is a range of alignable entities $A_{1}\subset\mathcal{E}_{1}$ and $A_{2}\subset\mathcal{E}_{2}$ (for DBP15K, $A_{1}\subsetneqq\mathcal{E}_{1}$). Second, the equivalence between $A_{1}$ and $A_{2}$ is a bijection. Note that the assumption does not have aligning regularity for entities outside the range except that they can’t be aligned with entities inside the range. Many ranking-based EA methods leverages the 1-to-1 range assumption, however, NALA do not. Therefore, in implementation in order to leverage the range assumption we take the set $A_{1}$ and $A_{2}$ as input and filters out any alignment sentence that aligns $A_{1}$ to $\mathcal{E}_{2}\setminus A_{2}$ or $\mathcal{E}_{1}\setminus A_{1}$ to $A_{2}$.
D.4 Relation with Other Methods
The way NAL models KG information and the inference process has a similar part with "uncertainty estimation"  Hu et al. (2023) in the natural language processing domain. The truth-value of alignments shares some similarity with the distributive view of facts or beliefs which views facts as probability distribution of random variables. Also, the concept of confidence is shared with some information extraction systems such as Markov logic network  Jiang et al. (2012), which assigns confidence to extracted facts or logical formulas in some intermediate steps.
D.5 Relation with Translational Embedding Methods
The well-known KG embedding model TransE Bordes et al. (2013) is initially proposed for link prediction tasks. It may be partially explained from a logical perspective of NAL (or equivalently other logic with similar expressive power). Consider a specific type of Horn clauses $((*,A,B)\rightarrow R_{1}\wedge\ (*,B,C)\rightarrow R_{2})\ \Rightarrow(*,A,C)\rightarrow R_{3}\ \left\langle f_{1},c_{1}\right\rangle$, the following three triples
$(Georgia\_(U.S.\_state),\ country,\ United\_States)$
$(Martin\_Luther\_King\_Jr,\ citizenship,\ United\_States)$
[…]
InCountry
manufacturercountrymade-
InCountry
[…]
. The training process may force vector birthPlace to be dissimilar with manufacturer, otherwise, there may be hallucination in link prediction or EA results. A similar explanation of hallucination may apply to LLMs. A similar analysis applies to the vector representations of two relations which frequently appear on the same head entity (or tail entity). It’s arguable that the test set link prediction process of TransE mainly relies on Horn clauses, because from a logical perspective there is scarcely any other information. In this paper Horn clauses will not be extracted and managed, leaving for further research.
MTransE  Chen et al. (2016) is a translational embedding-based EA method. It encodes the two KGs’ relational triples separately with the TransE loss criterion $S_{K}=\Sigma_{(h,r,t)}||\textbf{h}+\textbf{r}-\textbf{t}||$. It proposed a "distance-based axis calibration" alignment model in order to coincide the vectors of counterpart entities or relations. The corresponding loss is $S_{a_{2}}=\Sigma||\textbf{e}_{1}-\textbf{e}_{2}||+||\textbf{r}_{1}-\textbf{r}_{2}||$ ($S_{a_{2}}$ only has the first item if there is no available seed relation alignment). The seed and derived alignments are assumed to have $\textbf{e}_{1}\approx\textbf{e}_{2}$ and we see it as the embedding representation of the similarity statement $e_{1}\leftrightarrow e_{2}$, with its truth-value somehow represented by the distance $||\textbf{e}_{1}-\textbf{e}_{2}||$. Theoretically, the distance can’t simultaneously represent frequency and confidence by itself, but more possibly a combined effect. We argue that MTransE performs approximate inference that is similar with the type III path, because if the learned embedding constraints of the four premises are considered simultaneously, we can get $\textbf{r}_{1}\approx\textbf{r}_{2}$ which we interpret as $r_{1}\leftrightarrow r_{2}$. Similarly, MTransE performs approximate inference of the type I path (with functionality omitted and $r_{1}\rightarrow r_{2}$ replaced by $r_{1}\leftrightarrow r_{2}$) to obtain derived alignment results.
D.6 Relation with Embedding-path EA Methods
The first method to be discussed is BERT-INT. It generates entity embedding using the name/description information with BERT unit and the embedding is $C(e)=MLP(CLS(e))$. It uses pairwise margin loss to approximately enforce $C(e)\approx C(e^{\prime})$. Different from MTransE which performs path inference implicitly with the gradient optimization of loss criterions, BERT-INT explicitly performs path inference with its proposed interaction model. Every element of the neighbor-view interaction matrix represents a inference process of a type I path. Its path omits functionality and relation alignment (for BERT-INT fails to utilize its proposed relation mask matrix). Because of the ignorance of relation type, its premise (1) and (4) has the form of $\left(*,x_{1},y_{1}\right)\rightarrow\#r$ and $\left(*,x_{2},y_{2}\right)\rightarrow\#r$ which represents "There exists an unspecified relation between $x_{1}$/$y_{1}$, and (another) unspecified relation between $x_{2}$/$y_{2}$". Moreover, its premise (5) fails to utilize derived alignments, because BERT-INT is not iterative. With such premises, BERT-INT’s type I path inference’s effectiveness is supposed to be lower than that of NALA’s. Similarly, every element of the attribute-view interaction matrix represents a type I path which has attribute triples as premises (1) and (4). BERT-INT’s evidence aggregation method is different from NALA which uses probabilistic revision and revision rules.
In this paper, BERT-INT and FGWEA are classified as embedding-path EA methods because their embedding module couples with the path inference to some extent. In contrast, NALA, which we classify as path-based, performs path inference wherever it can and uses embeddings minimally.
Appendix E Experiments
Dataset $||\mathcal{E}|$$||\mathcal{R}|$$|\mathcal{T_{R}}|$ $|\mathcal{T_{A}}|$
$DBP15K_{ZH\_EN}$ 19,388 1,701 70,414 379,684
19,572 1,323 95,142 567,755
$DBP15K_{JA\_EN}$ 19,814 1,299 77,214 354,619
19,780 1,153 93,484 497,230
$DBP15K_{FR\_EN}$ 19,661 903 105,998 528,665
19,993 1,208 115,722 576,543
$DBP15Kfull_{ZH\_EN}$ 66,469 2,830 153,929 379,684
98,125 2,317 237,674 567,755
$DBP15Kfull_{JA\_EN}$ 65,744 2,043 164,373 354,619
95,680 2,096 233,319 497,230
Table 3: Dataset statistics. $|\mathcal{E}|$, $|\mathcal{R}|$, $|\mathcal{T_{R}}|$ and $|\mathcal{T_{A}}|$ represent the number of entities, relation types, relation triples and attribute triples in each KG, respectively. The statistics of OpenEA benchmark datasets are in  Sun et al. (2020)
E.1 Evaluation Metric & Environment
We use Hits@1 (which is the same metric as recall for EA) as the sole evaluation metric of our main results of DBP15K for the following reasons. Mean Reciprocal Rank (MRR) is unavailable for NALA because it does not provide a alignment ranking for the test entities. There exist a non-negligible number of equivalent entity pairs that are not in the ground-truth set of DBP15K, so the precision and F1-score can’t be measured properly. We use the precision (P), recall (R), and F1 score in the ablation study of OpenEA benchmark datasets.
Our NALA model is implemented in java and the BERT unit is implemented in python with PyTorch. All experiments are performed on a Linux server with an Intel(R) Xeon(R) Silver 4210R CPU @ 2.40GHz, 251G RAM and a NVIDIA GeForce RTX 3090 GPU.
E.2 Experimental Results on Full Version of DBP15K
The aforementioned and widely used DBP15K dataset version is reduced from a full version. That is, any triple that does not contain an entity in the 15k-15k 1-to-1 range is discarded, resulting in  19k entities in both side of KG. Thus, the full version has more entities and relation triples. Most embedding-based EA methods rely on the 1-to-1 range information to generate the rankings for alignment. For the full version of DBP15K, such information may have a bigger effect on EA performance. As shown in Table 4, to our knowledge only few methods have conducted experiments on the full version, yet none of them is implemented without 1-to-1 range. The results of other methods are from  Cai et al. (2022) (RPR-RHGT). NALA has reasonable Hits@1 performance even without the help of 1-to-1 range information and greatly surpasses other listed methods.
Method $full_{ZH\_EN}$ $full_{JA\_EN}$
NALA(without 1-to-1 range) 0.908 0.889
JAPE(with 1-to-1 range) 0.264 0.238
RDGCN(with 1-to-1 range) 0.621 0.812
RPR-RHGT(with 1-to-1 range) 0.693 0.886
NALA(with 1-to-1 range) 0.966 0.946
Table 4: Experimental results (Hits@1) on full version of DBP15K. The experiments for NALA are based on setting group 1, without attribute value embeddings.
E.3 Influence of Confidence Hyper-parameter
The experiment results of Figure 6 shows how entity name/description embedding similarity confidence $C_{name}$ (without the adaptive setting of $C_{name}$) affects Hits@1. These experiments are performed on setting group 4 without using attribute value embedding information. We adjust $C_{name}$ with other conditions unchanged. The Hits@1 curve is approximately concave and for $ZH\_EN$, $JA\_EN$ and $FR\_EN$ respectively, it reaches maximum performance at 0.6, 0.55 and 0.8. It shows that the informative embedding similarity enhances the performance to different extents. French is often regarded as more closely related to English than Chinese or Japanese, so the BERT unit learns representation easier and thus produces more confident embedding similarity. Pretraining corpus of the BERT unit may include relevant triples (in the form of natural language sentences) which may have same informational origin with DBpedia. So the embedding similarity’s evidences may have an overlap part with type I path’s evidences. The revision rule is only appropriately used when the two premises don’t share same evidence (or equivalently their evidential bases do not overlap). So the appropriate confidence value need to be lower than the confidence of the BERT output (if it provides such information) in order to exclude the overlap. The best-performance confidence of each dataset is conjectured to reflect the combined influence of embedding quality of the BERT unit and the evidence overlapping effect. The $C_{name}$ confidence value can be alternatively set equal to the cosine similarity of the embeddings, resulting in a slightly decreased performance. This is another good choice if you want to avoid hyper-parameter tuning.
Figure 6: Influence of $C_{name}$.
Dataset LightEA-I NALA NALA(w/o attr)
D-W-15K-V1 0.732 0.780 0.625
D-W-100K-V1 0.642 0.732 0.528
D-Y-15K-V1 0.826 0.949 0.751
D-Y-100K-V1 0.781 0.927 0.700
D-W-15K-V2 0.951 0.938 0.902
D-W-100K-V2 0.926 0.919 0.857
D-Y-15K-V2 0.976 0.983 0.975
D-Y-100K-V2 0.977 0.984 0.913
Table 5: Experimental results (Hits@1) on OpenEA benchmarks. The experiments for NALA are done without attribute value embeddings.
Figure 7: Results of bootstrap steps of setting group 3.
Figure 8: Results of the iterations of the second bootstrap step in setting group 3.
Appendix F Algorithms
@@
4       match_and_delete($e_{1},null$);
@@
/* $e_{1}$ is the entity to be matched and we assume that $e_{1}$ belongs to $\mathcal{KG}_{1}$, similarly otherwise. Entity $e_{prev}$ represents the previous entity, that is the concerned entity of the recursion parent. */
output : entity $e_{match}$ which forms a stable matching with $e_{1}$
@@ 3       $e_{2}$leftarrow $predicate\_term$ of $sentence$;
5             $e_{match}$leftarrow $e_{prev}$;
@@
9             $e_{3}$leftarrow match_and_delete($e_{2},e_{1}$);
@@ 11                   $e_{match}$leftarrow $e_{2}$;
Algorithm 2 match and delete
@@
8             for $x_{1}$, $x_{2}$, $y_{2}$ that forms a sound type I path with $y_{1}$ (depth-first) do
9                   perform inference of type I path;
10                   perform inference of type III path;
11
@@
14                   perform inference of type II path;
15